OpenStudy (anonymous):

sqrt(10^2.6 x 10^.5 / 10^1.5) my math book says leave the answer in radical or power form, not a decimal. I keep getting sqrt(10^8/5) but I'm unsure.

OpenStudy (anonymous):

\[2.6+.5-1.5=1.6=\frac{8}{5}\] \[\sqrt{10^{\frac{8}{5}}}\] \[10^{\frac{8}{10}}=10^{\frac{4}{5}}\]

OpenStudy (anonymous):

Not sure why you said 8/10 = 4/5, since it's 8/5, but thanks for confirming my answer.